How would I figure out how long it will take a principal to grow? How long would it take a $250 investment with a compounding APY of 1% to reach $300?
Answer by Alan3354(69443)

How would I figure out how long it will take a principal to grow? How long would it take a $250 investment with a compounding APY of 1% to reach $300?
------------------
You have to know the compounding period.
If it's annual, --> FV = PV*(1.01)^t
300 = 250*1.01^t
1.01^t = 1.2
t*log(1.01) = log(1.2)
t = log(1.2)/log(1.01) years
t =~ 18.32 years
=================
If it's monthly:
300 = 250*(1 + 0.01/12)^(t*12) --- 1/12 the interest, 12 times as many periods
1.2 = (1 + 0.01/12)^m --- m in months
m = log(1.2)/log(1 + 0.01/12)
m =~ 218.9 months
=~ 18.24 years